# Post-Newtonian templates for phase evolution of spherical extreme mass ratio inspirals

### v2

@@ IV.4 Performance of the frequency-domain Taylor approximants
In Figs. 13 and 14, we show the mismatch as a function of observation time. We also show the line with $\mathcal{M}_{\textrm{F2}}^{(n)}=10^{-3}$. This corresponds to the limit of mismatch so that two waveforms are indistinguishable, which is estimated by $\mathcal{M}\sim 1/(2\rho^{2})$ with the SNR of $\rho\simeq 20$, and roughly corresponds to the dephasing of $\Delta\Phi\sim O(0.1)$ in time-domain templates [34, 35, 39]. For System1 (early stage of inspiral), the 8PN order approximant keeps the mismatch less than $O(0.01)$ for two-year observation, but it is not sufficient to suppress the mismatch to $O(10^{-3})$. The 9th or higher PN order approximants are required. The performance of the TaylorF2 approximant is similar to the TaylorT2 in the time-domain families. As shown in Sec. IV.3, the 9th or higher PN TaylorT2 approximants are required to suppress the dephasing to $O(0.1)$ (corresponding to a $O(10^{-3})$ mismatch in the frequency domain) for two years. This fact may be expected from the similarity in the formulation of the TaylorT2 and TaylorF2 approximants.
For System2 (late stage of inspiral), even the 12PN approximants, which is the highest order we have currently, does not keep the mismatch less than $O(10^{-3})$ for two years. This tendency is same as we show for the phase evolution in the TaylorT families.
@@ V Summary
In this work, we presented various PN template families for the phase evolution of quasi-spherical EMRIs in Kerr spacetime and examine the convergence by evaluating the dephasing between approximants with different PN orders belonging to each family. We found that the PN convergence slightly depends on the black hole spin, $q$, and the initial value of the inclination, $Y_{I}$, and that the convergence gets slow when $q$ or $Y_{I}$ becomes large. This tendency implies that the rotation of the central black hole works to slow the convergence. In addition, the tendency suggests that we may focus on the cases with larger $q$ and larger $Y_{I}$ (the case of $Y_{I}=0.9$ and $q=0.9$ in the current work) to discuss the convergence of various approximants.
From the comparison of TaylorT families, we found that the TaylorT1 approximant shows the best performance among them as shown for circular orbits in Schwarzschild case [29]. For early inspirals (represented by System1), the 8PN TaylorT1 template is expected to keep the dephasing less than $O(0.1)$ during two-year observation. The alternative approximant of the TaylorT1, named TaylorT1a in this work, shows the equivalent performance too. As for the TaylorT2 template, the corrections at least up to the 9PN order will be required to achieve the comparable accuracy to the TaylorT1. Although the convergence of the TaylorT2 approximant is slightly slower than that of the TaylorT1/T1a, however, the fully analytical expression is useful to reduce the computational cost to construct the bank of GW templates and to discuss effects beyond general relativity.
To examine the convergence of the TaylorF2 approximant, we evaluated the mismatch between templates with different PN orders. For early inspirals, the corrections at least up to 9PN order will be required to keep the mismatch to $O(10^{-3})$, which corresponds to the dephasing of $O(0.1)$ in the time domain. This results supports the expectation that the TaylorF2 approximant has comparable performance to the TaylorT2 from the similarity in the formulation. Since the TaylorF2 approximant is given in fully analytical expression same as the TaylorT2, it provides an easy-to-handle method of calculating the frequency-domain templates for EMRIs.

### v3

@@ IV.3 Performance of the frequency-domain Taylor approximants
In Figs. 11 and 12, we show the mismatch as a function of observation time. We also show the line with $\mathcal{M}_{\textrm{F2}}^{(n)}=10^{-3}$. This corresponds to the limit of mismatch so that two waveforms are indistinguishable. This is estimated by $\mathcal{M}\sim 1/(2\rho^{2})$ with the SNR of $\rho\simeq 20$, and roughly corresponds to the dephasing of $\Delta\Phi\sim O(0.1)$ in time-domain templates [42, 43, 47]. For System1 (early stage of inspiral), the 8PN order approximant keeps the mismatch less than $O(0.01)$ for two-year observation, but it is not sufficient to suppress the mismatch to $O(10^{-3})$. The 9th or higher PN order approximants are required. The performance of the TaylorF2 approximant is similar to the TaylorT2 in the time-domain families. As shown in Sec. IV.2, the 9th or higher PN TaylorT2 approximants are required to suppress the dephasing to $O(0.1)$ (corresponding to a $O(10^{-3})$ mismatch in the frequency domain) for two years. This fact may be expected from the similarity in the formulation of the TaylorT2 and TaylorF2 approximants.
For System2 (late stage of inspiral), even the 12PN approximants, which is the highest order we have currently, does not keep the mismatch less than $O(10^{-3})$ for two years. This tendency is the same as we show for the phase evolution in the TaylorT families.
@@ V Summary
In this work, we presented various PN template families for the phase evolution of quasi-spherical EMRIs in Kerr spacetime, and examined the convergence by evaluating the dephasing between approximants with different PN orders belonging to each family. We found that the PN convergence slightly depends on the black hole spin, $q$, and the initial value of the inclination, $Y_{I}$, and that the convergence gets slow when $q$ or $Y_{I}$ becomes large. This tendency implies that the rotation of the central black hole works to slow the convergence. In addition, the tendency suggests that we may focus on the cases with larger $q$ and larger $Y_{I}$ (the case of $Y_{I}=0.9$ and $q=0.9$ in the current work) to discuss the convergence of various approximants.
From the comparison of TaylorT families, we found that the TaylorT1 approximant shows the best performance among them as shown for circular orbits in Schwarzschild spacetime [38]. For early inspirals (represented by System1), the 8PN TaylorT1 template is expected to keep the dephasing less than $O(0.1)$ during two-year observation. The alternative approximant of the TaylorT1, named TaylorT1a in this work, shows the equivalent performance too. As for the TaylorT2 template, the corrections at least up to the 9PN order will be required to achieve the comparable accuracy to the TaylorT1. The convergence of the TaylorT2 approximant is slightly slower than that of the TaylorT1/T1a. But the fully analytical expression is useful to reduce the computational cost to construct the bank of GW templates and to discuss effects beyond general relativity.
To examine the convergence of the TaylorF2 approximant, we evaluated the mismatch between templates with different PN orders. For early inspirals, the corrections at least up to 9PN order will be required to keep the mismatch to $O(10^{-3})$, which corresponds to the dephasing of $O(0.1)$ in the time domain. This result supports the expectation that the TaylorF2 approximant has comparable performance to the TaylorT2 from the similarity in the formulation. Since the TaylorF2 approximant is given in fully analytical expression same as the TaylorT2, it provides an easy-to-handle method of calculating the frequency-domain templates for EMRIs.
@@ Appendix H Expansion coefficients of $\mathcal{E}_{n}(x)$ and $\mathcal{F}_{n}(x)$
Appendix J Convergence of PN formulas for $E$ and $L$
In a similar manner to Eq. (67), we check the convergence of the PN formulas of the specific energy and angular momentum, Eqs. (20) and (21), by introducing the difference
In Figs. 13 and 14, we show $\Delta E^{(n)}$ and $\Delta L^{(n)}$ as functions of $x$ for several sets of $(q,Y)$. In these plots, $\Delta E^{(n)}$ and $\Delta L^{(n)}$ are normalized by the leading (Newtonian) order terms,
$E_{N}\equiv 1-\frac{1}{2}x^{2},\quad L_{N}\equiv\frac{Y}{x}.$
We find a weak dependence on $q$ in Figs. 13 and 14: the PN convergence gets slightly worse when $q$ increases. On the other hand, we find no dependence on $Y$ in these plots at least with one’s eyes.
Figure 13: Contribution of each correction term in the PN expansion of the specific energy. We plot $\Delta E^{(n)}$ normalized by the Newtonian energy, $E_{N}$, as functions of $x$ for several sets of $(q,Y)$. In the same reason as Fig. 1, we show only the 2PN, 4PN, 6PN, 8PN, 10PN and 12PN corrections in this figure.